To whom may it concern,
I have problem to do this question.
The heights of three plants A, B and C in a garden are in the ratio of 2:3:5. Their mean height is 30 cm.
(a) Find the height of plant B
(b) If another plant D is added to the garden and the mean height of the four plants is now 33 cm, find the height of plant D.
Please help. Thank you so much.
CH Lam

Answer by stanbon(48568)

The heights of three plants A, B and C in a garden are in the ratio of 2:3:5. Their mean height is 30 cm.
(a) Find the height of plant B
(b) If another plant D is added to the garden and the mean height of the four plants is now 33 cm, find the height of plant D.
----------
Heights are:
A : 2x
B : 3x
C ; 5x
Average height is 30 cm
Add them up and divide by 3 to get 30.
(2x+3x+5x)/3 = 30
Solve for "x".
10x=90
x=9
Determine B's height:
B=3x=3(9)= 27 cm
------
Figure the height of A and C
A=2x=18
C=5x=45
Add A,B,C, and D and divide by 4 to get 33
(18+27+45+D)/4=33
90+D=132
D=42 cm.
Cheers,
Stan H.